In a certain region of space electric field is along the z direction throughout.the magnitude of electric field is however not constant but increase uniformly along the positive direction of z axis.at the rate of 10*5 N/C per metre.what are the force and torque experienced by a system having a total dipole moment of 10*-7 Cm in the negative z direction

Grade:12

1 Answers

Eshan
5 years ago
Dear student,

Since the dipole moment and the electric field, both are along z-axis, no torque acts on the body.
Let the magnitude of charges beqand the distance between the charges bed.
Hence net force acting on the dipole=q_1E_1+q_2E_2=qE_1-qE_2=q(E_1-E_2)=q(\dfrac{\Delta E}{\Delta z}d)=qd(\dfrac{\Delta E}{\Delta z}))=10^{-7}\times 10^5N=0.01N